OpenStudy (anonymous):

Graph and check to solve the system of linear equations. 2x + y = 4 -4x – 2y = -8

OpenStudy (anonymous):

need help

OpenStudy (texaschic101):

I know how to solve it but I am no good at graphing. Do you want me to show you how to solve it without graphing ?

OpenStudy (anonymous):

i just need to solve it

OpenStudy (texaschic101):

do you know how to do the elimination method ?

OpenStudy (anonymous):

not really

OpenStudy (texaschic101):

It can also be solved by substitution....that is easier, so I will show you that... 2x + y = 4 (you want to get the single variable, y, on one side by itself. so subtract 2x from both sides.. 2x - 2x + y = 4 - 2x (combine like terms) y = 4 - 2x Do you understand how I did this ?

OpenStudy (anonymous):

yes but how do you get the points

OpenStudy (texaschic101):

I will show you....be patient. we now know that y = 4 - 2x. substitute 4 - 2x in for y in the other equation.. -4x - 2y = -8 -4x -2(4 - 2x) = -8 (distribute the -2 through the parenthesis) -4x - 8 + 4x = - 8 -4x +4x = -8 + 8 0 = 0 There is infinite solutions to this problem.

OpenStudy (anonymous):

thanks i see how u did it

OpenStudy (texaschic101):

if the answer would have been 0 = 3, then there would be no solutions

hero (hero):

A simpler way to do it would have been to just take the second equation: -4x – 2y = -8 Rewrite it as: 4x + 2y = 8 Then divide by 2 to get: 2x + y = 4 Thus both equations are the same

OpenStudy (texaschic101):

yes....that is much easier :)